OpenStudy (rissaunnie):

does anyne know of a calculator i can use for my Translations and Transformations homework

OpenStudy (kendricklamar2014):

Does this site help?: https://www.desmos.com/calculator/c2kkis2aog

OpenStudy (rissaunnie):

no it tells me invalid input

OpenStudy (kendricklamar2014):

Hmmm.. http://www.idomaths.com/linear_transformation.php

OpenStudy (rissaunnie):

nope]

OpenStudy (kendricklamar2014):

Hmm thats all the calculators I know. Sorry for the "help" :(

OpenStudy (rissaunnie):

thanks for trying cx

OpenStudy (aric200):

Can you give me a sample question ?

OpenStudy (rissaunnie):

suppose f(x)=\[x^{3}\] find the graph of f(x+3)

OpenStudy (mrnood):

you don't need a calculator - you need read your text and undersatand the question calculators are for finding the answer WHEN you know how to do it. They do not normally have a 'do homework' button in my experience. the funnction takes the 'input' of x and give x^3 as the out put so - before oyu try to 'find the graph' you need to write down the function as it will look when the 'input' "x+3" FOR instance if f(x) = x^2 then to find f(x-1) we can see f(x-1) = (x-1)^2 this is similar to your question - so use that method to write f(x+3)

OpenStudy (anonymous):

What about mathway?